# Multi-Armed Bandits with Abstention

### 23 Feb 2024 v1

@@ Appendix A Asymptotic and Minimax Optimality in Canonical Multi-Armed Bandits
Definition A.2 ($R^{\mathrm{CA}}$-consistency).
@@
According to Theorem A.3, in the canonical bandit model, an algorithm $\pi\in\Pi^{\mathrm{CA}}$ is said to be asymptotically optimal if for all bandit instances $\mu\in\mathcal{U}$, it ensures that
@@
[…]
Remark A.4.
One valuable byproduct derived from the proof of the asymptotic lower bound in Theorem A.3 is that, for any $R^{\mathrm{CA}}$-consistent algorithm $\pi$, bandit instance $\mu\in\mathcal{U}$ and suboptimal arm $i>1$, we have
@@
Appendix B Auxiliary Lemmas
Lemma B.1 (Bretagnolle–Huber inequality (Tsybakov, 2009)).
@@ Appendix B Auxiliary Lemmas
where KL(,)normal-$\mathrm{KL}(\mathbb{P},\mathbb{P}^{\prime})$ denotes the Kullback–Leibler (KL) divergence between $\mathbb{P}$ and $\mathbb{P}^{\prime}$.
Lemma B.2 (Divergence decomposition lemma).
@@
The proof of Lemma B.2 is similar to the well-known proof of divergence decomposition in the canonical multi-armed bandit model (excluding abstention), and is therefore omitted. This proof can be located, for instance, in Garivier et al. (2019, Section 2.1) and Lattimore & Szepesvári (2020, Lemma 15.1).
Lemma B.3 (Hoeffding’s inequality for sub-Gaussian random variables).
@@
$\mathbb{P}(\hat{\mu}\geq\mu+\varepsilon)\leq\exp\left(-\frac{n\varepsilon^{2}}{2\sigma^{2}}\right)\quad\text{ and }\quad\mathbb{P}(\hat{\mu}\leq\mu-\varepsilon)\leq\exp\left(-\frac{n\varepsilon^{2}}{2\sigma^{2}}\right)$
@@
$\sum_{n=1}^{N}\mathbb{P}(\hat{\mu}_{n}\geq\mu+\varepsilon)\leq\frac{2\sigma^{2}}{\varepsilon^{2}}\quad\text{ and }\quad\sum_{n=1}^{N}\mathbb{P}(\hat{\mu}_{n}\leq\mu-\varepsilon)\leq\frac{2\sigma^{2}}{\varepsilon^{2}}$
@@
Proof of Lemma B.4.
By symmetry, it suffices to prove the first part. According to Lemma B.3, we have
@@
Appendix C Analysis of the Fixed-Regret Setting
C.1 Upper Bounds
Proof of Theorem 3.3.
@@ C.1 Upper Bounds
$R^{\mathrm{RG}}_{\mu,c}(T,\pi)$ $=\mathbb{E}\left[\sum_{t=1}^{T}\big{(}(\mu_{1}-X_{t})\cdot\mathbbm{1}\{B_{t}=0\}+c\cdot\mathbbm{1}\{B_{t}=1\}\big{)}\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}\big{(}(\mu_{1}-\mu_{A_{t}})\cdot\mathbbm{1}\{B_{t}=0\}+c\cdot\mathbbm{1}\{B_{t}=1\}\big{)}\right]$
@@
For convenience, for any $i\in[K]$ such that $\Delta_{i}<c$, we introduce three shorthand notations to represent the terms in (5):
$\begin{cases}\vspace*{4pt}(\clubsuit)_{i}:=\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\mathbb{I}\left\{A_{t}=i\text{ and }\sqrt{\frac{K}{t}}\geq c\right\}\right]\\ \vspace*{4pt}(\spadesuit)_{i}:=\operatorname{\mathbb{E}}\left[\sum_{t=K+1}^{T}\mathbb{I}\left\{A_{t}=i\text{ and }\max_{j\in[K]}\left(\hat{\mu}_{j}(t-1)-\sqrt{\frac{6\log t+2\log(c\vee 1)}{N_{j}(t-1)}}\right)\geq\mu_{1}\right\}\right]\\ (\blacksquare)_{i}:=\operatorname{\mathbb{E}}\left[\sum_{t=K+1}^{T}\mathbb{I}\left\{A_{t}=i\text{ and }\mu_{1}-\hat{\mu}_{i}(t-1)\geq c\right\}\right].\end{cases}$
We will deal with $(\clubsuit)_{i}$ and $(\spadesuit)_{i}$ later for the two forms of upper bounds.
On the other hand, for the term $(\blacksquare)_{i}$, we have
@@
For all $j\in[K]$, by a union bound over all possible values of $N_{j}(t-1)$ and Lemma B.3, we have
@@
where the term $\beta(b,\mu,K)$ is subsequently defined in Lemma C.1.
@@
Proof of Lemma C.1.
The proof of Lemma C.1 closely follows that of Proposition 5 in Korda et al. (2013), which was used to analyze the classical Thompson Sampling algorithm. In fact, the only difference between our arm sampling rule, which is built upon Less-Exploring Thompson Sampling (Jin et al., 2023), and the classical Thompson Sampling is how the estimated reward $a_{i}(t)$ is constructed for each arm $i\in[K]$. Specifically, in our arm sampling rule, $a_{i}(t)$ is either drawn from the posterior distribution $\mathcal{N}(\hat{\mu}_{i}(t-1),1/N_{i}(t-1))$ with probability $1/K$ or set to be the empirical mean $\hat{\mu}_{i}(t-1)$ otherwise. In classical Thompson Sampling, $a_{i}(t)$ is always drawn from the posterior distribution. Therefore, it suffices to verify the parts concerning the probability distributions of $a_{i}(t)$; these correspond to Lemmas 9 and 10 in the proof of Proposition 5 in Korda et al. (2013).
It is straightforward to see that Lemma 9 in Korda et al. (2013) is applicable to our algorithm. For Lemma 10 therein, its counterpart is demonstrated in Lemma C.2 below.
@@
  • First, for a suboptimal arm, if it has been pulled a sufficient number of times, then, with high probability, its estimated reward (sample) cannot deviate significantly from its true mean. This observation is quantitatively characterized in Lemma 10 of Korda et al. (2013), corresponding to Lemma C.2 in our paper.
@@
For the first term in (17), by Lemma B.3, we can bound it as
@@
For the second term in (17), according to the construction of $a_{i}(s)$ in Algorithm 1 and Lemma B.3, we can get
@@
C.2 Lower Bounds
Proof of Theorem 3.5.
@@ C.2 Lower Bounds
Therefore, in accordance with Definition A.2, the algorithm $\pi$ is also $R^{\mathrm{CA}}$-consistent for arbitrary abstention regret $c$.
@@
where the last inequality follows from the property of $R^{\mathrm{CA}}$-consistent policies as detailed in Remark A.4.
@@
By combining the aforementioned two inequalities and applying Lemma B.1, we obtain the following:
@@
Leveraging Lemma B.2 and the KL divergence between Gaussian distributions, we can derive
@@
Appendix D Analysis of the Fixed-Reward Setting
D.1 Upper Bounds
Proof of Theorem 4.1.

### 22 Mar 2026 v2

@@ Appendix A Asymptotic and Minimax Optimality in Canonical Multi-Armed Bandits
Definition 3 ($R^{\mathrm{CA}}$-consistency).
@@
According to Theorem 5, in the canonical bandit model, an algorithm $\pi\in\Pi^{\mathrm{CA}}$ is said to be asymptotically optimal if for all bandit instances $\mu\in\mathcal{U}$, it ensures that
@@
To the best of our knowledge, for canonical multi-armed bandits with Gaussian rewards, $\text{KL-UCB}^{++}$ (Ménard & Garivier, 2017), ADA-UCB (Lattimore, 2018), MOTS-$\mathcal{J}$ (Jin et al., 2021) and Less-Exploring Thompson Sampling (Jin et al., 2023) exhibit simultaneous asymptotic and minimax optimality. As their names suggest, the former two algorithms follow the UCB-style, while the latter two are rooted in Thompson Sampling.
Remark 8.
One valuable byproduct derived from the proof of the asymptotic lower bound in Theorem 5 is that, for any $R^{\mathrm{CA}}$-consistent algorithm $\pi$, bandit instance $\mu\in\mathcal{U}$ and suboptimal arm $i>1$, we have
@@
Appendix B Comparison with the $(K+1)$-Armed Bandit Reduction
A natural question arises regarding the formulation of our abstention model: can the problem be simply reduced to a standard multi-armed bandit setting with $K+1$ arms? In such a reduction, the additional $(K+1)$-st arm would represent the abstention option, yielding a deterministic reward $c$ (in the fixed-reward setting) or a fixed regret (in the fixed-regret setting).
While such a reduction might appear intuitive and indeed may suffice for achieving coarse worst-case (minimax) guarantees of $O(\sqrt{KT})$, it fundamentally fails to capture the strategic advantages of our framework. Specifically, the reduction cannot achieve the same instance-dependent asymptotic optimality. In the following, we discuss the essential differences between our model and the $(K+1)$-armed reduction.
Structural Difference: Decoupling Learning from Reward.
The most critical distinction lies in the mechanism of information acquisition. In a standard $(K+1)$-armed bandit model, selecting the deterministic $(K+1)$-th arm (the “safe” option) fixes the instantaneous outcome but reveals no information regarding the distributions of the original $K$ stochastic arms. To learn about a specific arm $i\in[K]$, the agent is forced to pull it and fully accept its stochastic outcome.
In contrast, our abstention framework allows the agent to decouple the observation of an outcome from the acceptance of its reward. When the agent exercises the abstention option, they secure a guaranteed result (equivalent to the deterministic arm) but, crucially, simultaneously observe the stochastic sample $X_{t}$ from the selected arm $A_{t}$. This structure enables what can be termed “safe exploration”: the agent can continue to gather valuable statistics about potentially risky or unknown arms to refine their estimates, all while shielding themselves from the full cost of stochastic underperformance.
Performance Gap: Asymptotic Optimality.
To illustrate how this structural difference translates into superior asymptotic performance, we show below that in both the fixed-reward and fixed-regret objectives, the $(K+1)$-armed reduction and our abstention model admit different instance-dependent asymptotic behavior.
Fixed Reward: In a standard $(K+1)$-armed bandit reduction, to learn about a suboptimal arm $i$, the agent must select it. Consequently, the agent incurs a regret proportional to the full suboptimality gap $\Delta_{i}=\mu_{1}-\mu_{i}$ for every exploratory sample. The asymptotic regret contribution from arm $i$ is thus driven by the factor $\Delta_{i}$. However, in our fixed-reward abstention model, the agent can leverage the abstention option to mitigate this cost. Specifically, if the abstention reward $c$ satisfies $\mu_{i}<c<\mu_{1}$, the agent can pull arm $i$ while simultaneously choosing to abstain. In this scenario, the agent observes the outcome of arm $i$ (gaining information) but receives the fixed reward $c$. The resulting instantaneous regret is $\mu_{1}-c$, which is strictly less than the full gap $\mu_{1}-\mu_{i}$.
This mechanism facilitates a more cost-effective form of exploration. Consequently, our derived asymptotic regret bound scales with the term $\mu_{1}\vee c-\mu_{i}\vee c$, rather than the larger gap $\Delta_{i}$ that arises from the (K+1)-armed reduction. This confirms that a simple reduction to a $(K+1)$-armed bandit cannot replicate the superior asymptotic optimality achieved by our proposed model.
Fixed Regret: A simple $(K+1)$-armed reduction also fails in the fixed-regret objective. In that reduction, the $(K+1)$-st “abstain” arm incurs a deterministic regret $c$, but it reveals no information about the original $K$ stochastic arms. Hence, to statistically verify that a suboptimal arm $i$ is worse than the best arm, the learner must still pull arm $i$ and accept its stochastic outcome, paying regret about $\Delta_{i}$ per sample. By the standard asymptotic lower bound in Lai & Robbins (1985), the regret caused by pulling arm $i$ is on the order of $\frac{2\log T}{\Delta_{i}}$. Therefore, the leading asymptotic term under the $(K+1)$-arm reduction is $\sum_{i>1}\frac{2\log T}{\Delta_{i}}$, which is independent of $c$. In contrast, in our abstention model the learner can pull arm $i$ while abstaining, still observing $X_{t}$ but capping the instantaneous regret by $c$. Thus, exploration of arm $i$ costs $\Delta_{i}\wedge c$ per sample, yielding the strictly smaller scaling $2\sum_{i>1}\frac{\Delta_{i}\wedge c}{\Delta_{i}^{2}}\log T$, and in particular improves $\frac{2\log T}{\Delta_{i}}$ to $\frac{2c\log T}{\Delta_{i}^{2}}$ whenever $\Delta_{i}>c$. This again shows the (K+1)(K\!+\!1)-armed reduction cannot replicate the fixed-regret asymptotic achieved by our model.
Appendix C Auxiliary Lemmas
Lemma 1 (Bretagnolle–Huber inequality (Tsybakov, 2009)).
@@ Appendix C Auxiliary Lemmas
where $\mathrm{KL}(\mathbb{P},\mathbb{P}^{\prime})$ denotes the Kullback–Leibler (KL) divergence between $\mathbb{P}$ and $\mathbb{P}^{\prime}$.
Lemma 2 (Divergence decomposition lemma).
@@
The proof of Lemma 2 is similar to the well-known proof of divergence decomposition in the canonical multi-armed bandit model (excluding abstention), and is therefore omitted. This proof can be located, for instance, in Garivier et al. (2019, Section 2.1) and Lattimore & Szepesvári (2020, Lemma 15.1).
Lemma 3 (Hoeffding’s inequality for sub-Gaussian random variables).
@@
[…]
@@
[…]
@@
Proof.
By symmetry, it suffices to prove the first part. According to Lemma 3, we have
@@
Appendix D Analysis of the Fixed-Regret Setting
D.1 Upper Bounds
Proof of Theorem 1.
@@ D.1 Upper Bounds
$R^{\mathrm{RG}}_{\mu,c}(T,\pi)$ $=\mathbb{E}\left[\sum_{t=1}^{T}\big((\mu_{1}-X_{t})\cdot\mathbf{1}\{B_{t}=0\}+c\cdot\mathbf{1}\{B_{t}=1\}\big)\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}\big((\mu_{1}-\mu_{A_{t}})\cdot\mathbf{1}\{B_{t}=0\}+c\cdot\mathbf{1}\{B_{t}=1\}\big)\right]$
@@
For convenience, for any $i\in[K]$ such that $\Delta_{i}<c$, we introduce three shorthand notations to represent the terms in Equation (5). These terms correspond to distinct conditions under which the algorithm chooses to abstain:
$\begin{cases}\vskip 4.0pt(\clubsuit)_{i}:=\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\mathbb{I}\left\{A_{t}=i\text{ and }\sqrt{\frac{K}{t}}\geq c\right\}\right]\\ \vskip 4.0pt(\spadesuit)_{i}:=\operatorname{\mathbb{E}}\left[\sum_{t=K+1}^{T}\mathbb{I}\left\{A_{t}=i\text{ and }\max_{j\in[K]}\left(\hat{\mu}_{j}(t-1)-\sqrt{\frac{6\log t+2\log(c\vee 1)}{N_{j}(t-1)}}\right)\geq\mu_{1}\right\}\right]\\ (\blacksquare)_{i}:=\operatorname{\mathbb{E}}\left[\sum_{t=K+1}^{T}\mathbb{I}\left\{A_{t}=i\text{ and }\mu_{1}-\hat{\mu}_{i}(t-1)\geq c\right\}\right].\end{cases}$
To clarify the intuition behind these terms:
  • $(\clubsuit)_{i}$ represents the gap-independent abstention. This captures the case where the algorithm abstains purely because the time step $t$ is small, independent of the empirical estimates.
  • $(\spadesuit)_{i}$ represents abstentions caused by the overestimation of LCBs. This captures the low-probability event where the lower confidence bound of any arm $j$ exceeds the true mean of the optimal arm $\mu_{1}$.
  • $(\blacksquare)_{i}$ represents abstentions caused by the underestimation of the selected arm. This captures the event where the empirical mean of the chosen arm $i$ deviates significantly below its true mean $\mu_{i}$ (specifically such that $\mu_{1}-\hat{\mu}_{i}(t-1)\geq c$), triggering the gap-dependent abstention criterion.
We will analyze $(\clubsuit)_{i}$ and $(\spadesuit)_{i}$ later to establish the two forms of upper bounds.
On the other hand, for the term $(\blacksquare)_{i}$, which relates strictly to the concentration of arm $i$, we have
@@
For all $j\in[K]$, by a union bound over all possible values of $N_{j}(t-1)$ and Lemma 3, we have
@@
where the term $\beta(b,\mu,K)$ is subsequently defined in Lemma 5.
@@
Proof.
The proof of Lemma 5 closely follows that of Proposition 5 in Korda et al. (2013), which was used to analyze the classical Thompson Sampling algorithm. In fact, the only difference between our arm sampling rule, which is built upon Less-Exploring Thompson Sampling (Jin et al., 2023), and the classical Thompson Sampling is how the estimated reward $a_{i}(t)$ is constructed for each arm $i\in[K]$. Specifically, in our arm sampling rule, $a_{i}(t)$ is either drawn from the posterior distribution $\mathcal{N}(\hat{\mu}_{i}(t-1),1/N_{i}(t-1))$ with probability $1/K$ or set to be the empirical mean $\hat{\mu}_{i}(t-1)$ otherwise. In classical Thompson Sampling, $a_{i}(t)$ is always drawn from the posterior distribution. Therefore, it suffices to verify the parts concerning the probability distributions of $a_{i}(t)$; these correspond to Lemmas 9 and 10 in the proof of Proposition 5 in Korda et al. (2013).
It is straightforward to see that Lemma 9 in Korda et al. (2013) is applicable to our algorithm. For Lemma 10 therein, its counterpart is demonstrated in Lemma 6 below.
@@
  • First, for a suboptimal arm, if it has been pulled a sufficient number of times, then, with high probability, its estimated reward (sample) cannot deviate significantly from its true mean. This observation is quantitatively characterized in Lemma 10 of Korda et al. (2013), corresponding to Lemma 6 in our paper.
@@
For the first term in (17), by Lemma 3, we can bound it as
@@
For the second term in (17), according to the construction of $a_{i}(s)$ in Algorithm 1 and Lemma 3, we can get
@@
D.2 Lower Bounds
Proof of Theorem 2.
@@ D.2 Lower Bounds
Therefore, in accordance with Definition 3, the algorithm $\pi$ is also $R^{\mathrm{CA}}$-consistent for arbitrary abstention regret $c$.
@@
where the last inequality follows from the property of $R^{\mathrm{CA}}$-consistent policies as detailed in Remark 8.
@@
By combining the aforementioned two inequalities and applying Lemma 1, we obtain the following:
@@
Leveraging Lemma 2 and the KL divergence between Gaussian distributions, we can derive
@@
D.3 Heterogeneous Abstention Regret
In the scenario of heterogeneous abstention regret, for each arm $i\in[K]$, the agent incurs a regret of $c_{i}>0$ when pulling arm $i$ and choosing abstention. By the law of total expectation, the expected cumulative regret up to the time horizon $T$ can be expressed as:
$R^{\mathrm{RG}}_{\mu,\mathbf{c}}(T,\pi)$ $=\mathbb{E}\left[\sum_{t=1}^{T}\big((\mu_{1}-X_{t})\cdot\mathbf{1}\{B_{t}=0\}+c_{A_{t}}\cdot\mathbf{1}\{B_{t}=1\}\big)\right]$
$=c_{1}\cdot\operatorname{\mathbb{E}}[N_{1}^{(1)}(T)]+\sum_{i>1}\left(\Delta_{i}\cdot\operatorname{\mathbb{E}}[N_{i}^{(0)}(T)]+c_{i}\cdot\operatorname{\mathbb{E}}[N_{i}^{(1)}(T)]\right).$
For ease of presentation, we define $c=\max_{i\in[K]}c_{i}$ as the maximum abstention regret level. The lower bound part is straightforward. Specifically, we can obtain the following lower bounds.
Theorem 6.
For any abstention regret $\mathbf{c}=(c_{i})_{i\in[K]}$, bandit instance $\mu\in\mathcal{U}$ and $R^{\mathrm{RG}}$-consistent algorithm $\pi$, it holds that
$\liminf_{T\to\infty}\frac{R^{\mathrm{RG}}_{\mu,\mathbf{c}}(T,\pi)}{\log T}\geq 2\sum_{i>1}\frac{\Delta_{i}\wedge c_{i}}{\Delta_{i}^{2}}.$
For any maximum abstention regret level $c>0$ and time horizon $T\geq K$, there exists a universal constant $\alpha>0$ such that
$\inf_{\pi\in\Pi^{\mathrm{RG}}}\sup_{\mu\in\mathcal{U},\mathbf{c}}{R^{\mathrm{RG}}_{\mu,\mathbf{c}}(T,\pi)}\geq\alpha(\sqrt{KT}\wedge cT).$
With respect to the algorithm design, for each arm $i\in[K]$, we redefine its lower confidence bound as
$\mathrm{LCB}_{i}(t)=\hat{\mu}_{i}(t-1)-\sqrt{\frac{6\log t+2\log(c_{i}\vee 1)}{N_{i}(t-1)}}.$
Then in the abstention decision rule, we choose to abstain ($B_{t}=1$) if and only if
$\max_{i\in[K]\setminus\{A_{t}\}}\mathrm{LCB}_{i}(t)-\hat{\mu}_{A_{t}}(t-1)\geq c_{i}\ \text{ or }\ \sqrt{\frac{K}{t}}\geq c.$ (18)
Note that in (18), we compare $\sqrt{\frac{K}{t}}$ with $c=\max_{i\in[K]}c_{i}$.
Theorem 7.
For any abstention regret $\mathbf{c}=(c_{i})_{i\in[K]}$ and bandit instance $\mu\in\mathcal{U}$, Algorithm 1 with the new abstention criteria in (18) guarantees that
$\limsup_{T\to\infty}\frac{R^{\mathrm{RG}}_{\mu,\mathbf{c}}(T)}{\log T}\leq 2\sum_{i>1}\frac{\Delta_{i}\wedge c_{i}}{\Delta_{i}^{2}}.$
Furthermore, there exists a universal constant $\alpha>0$ such that
${R^{\mathrm{RG}}_{\mu,\mathbf{c}}(T)}\leq\begin{cases}cT&\text{ if }c\leq\sqrt{K/T}\\ \alpha(\sqrt{KT}+\sum_{i>1}{\Delta_{i}})&\text{ if }c>\sqrt{K/T}.\end{cases}$
The proof of Theorem 7 closely follows that of Theorem 1 in Appendix D.1. While many steps are analogous, it is noteworthy that in this context, for any arm $i$ with $\Delta_{i}<c_{i}$, we can upper bound $\operatorname{\mathbb{E}}[N_{i}^{(1)}(T)]$ as
$\operatorname{\mathbb{E}}[N_{i}^{(1)}(T)]$ $\leq\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\mathbb{I}\left\{A_{t}=i\text{ and }\sqrt{\frac{K}{t}}\geq c\right\}\right]$
$\phantom{\ =\ }+\operatorname{\mathbb{E}}\left[\sum_{t=K+1}^{T}\mathbb{I}\left\{A_{t}=i\text{ and }\max_{j\in[K]}\left(\hat{\mu}_{j}(t-1)-\sqrt{\frac{6\log t+2\log(c_{j}\vee 1)}{N_{j}(t-1)}}\right)\geq\mu_{1}\right\}\right]$
$\phantom{\ =\ }+\operatorname{\mathbb{E}}\left[\sum_{t=K+1}^{T}\mathbb{I}\left\{A_{t}=i\text{ and }\mu_{1}-\hat{\mu}_{i}(t-1)\geq c_{i}\right\}\right].$
Appendix E Analysis of the Fixed-Reward Setting
E.1 Upper Bounds
Proof of Theorem 3.